OpenStudy (anonymous):

2. Add and answer in simplest fractional form. 3/8 + -1/8 (Points : 1) 1/8 2/16 1/4 Question 3.3. Add and answer in simplest fractional form. -1/5 + -3/4 = (Points : 1) -4/9 -19/20 -4/20 Question 4.4. Add and answer in simplest fractional form. -4/3 + 7/8 = (Points : 1) -11/24 3/5 3/24 Question 5.5. If a rancher sold 1/3 of his cattle and then sold another 1/4 of the original amount, what part of his cattle was sold? (Points : 1) 1/7 2/12 7/12

OpenStudy (freckles):

\[\frac{3}{8}+\frac{-1}{8}=\frac{3+(-1)}{8}\] so since they had a common denominator you can write the fraction as one like i just did 3+(-1)=?